OpenStudy (anonymous):

3. The monopolist faces a demand curve given by D(p) = 10p-3 . Its cost function is c(y) = 2y. What is its optimal level of output and price?

OpenStudy (anonymous):

using diffrerential calculus

OpenStudy (anonymous):

-> R (y) = (10y -3) 2y = 20y² - 6y -> R' (y) = 40y - 6 Optimal level of output when R' (y) = C'(y) -> 40y - 6 = 2 => 40y = 8 => y = 8/40 = .2

OpenStudy (anonymous):

thanks but why did u substitute the p for y ? additionally i need to solve for p-price

OpenStudy (anonymous):

The reason why it takes so long because the question doesn't explain which one is quantity!

OpenStudy (anonymous):

oh my there is a correction also, the fuction is actually \[D(p)=10p ^{-3}\]

OpenStudy (anonymous):

The concept is find Revenue: quantity * unit price R (y ) = (10/ y³ ) * 2y = 20/ y² In order to gain optimal output: R' (x) = C'(x) --> - 20/ y = 2 => y = -10 Plug it back to get unit price!

OpenStudy (anonymous):

There's so much confused in the way the asker posting! I just show you the concept with sample, then you can figure out for next problems.

OpenStudy (anonymous):

thanks now i follow

OpenStudy (anonymous):

i think. still reading though, but kinda make sense

OpenStudy (anonymous):

optimal output = maximizie profit => marginal revenue = marginal cost!

OpenStudy (anonymous):

yea i know

OpenStudy (anonymous):

It's weird that the result negative ???

OpenStudy (anonymous):

What course is it? Do you have the examples to practice with?

OpenStudy (anonymous):

it Intermediate microeconomics II , by problem is the demand curve funtion, usually they give function int terps of quantity eg P=700 - 5Q but now the funtion is in terms of price so i was a bit confused, proably still am

OpenStudy (anonymous):

I'm reading the material, hopefully I'll be able to clarify it for you!

OpenStudy (anonymous):

thanks alot

OpenStudy (anonymous):

This is online example, both Price and Cost depends on quantity Q: P = 100 − Q ( price function from demand) C(Q) = 20 Q ( Cost function) Find the monopolist quantity (Qm), price (Pm): MR = MC 100 − 2Q = 20 =⇒ Qm = 40 units , Pm = 100 - 40 = $60

OpenStudy (anonymous):

i'm wondering where they got marginal revenue = 100 - 2Q

OpenStudy (anonymous):

I'm glad that I don't take this course, good luck!

OpenStudy (anonymous):

thanks, wish i didnt either :( lol

OpenStudy (anonymous):

im wondering if they took the q from the other side leaving the 20 and gettin 2q

OpenStudy (anonymous):

Hopefully you can find some useful from the link!